Information ABC Company Number of Employees 7 Employee Name Hours worked Hourly rate John Doe 40 $25.00 Jane Jones 40 $24.00 Jessie Smith 40 $17.00 Erik Ackers 25 $15.00 Nan Schmit 30 $17.00 Larry Johnson 35 $18.00 Christopher Hay 40 $18.00 Assume Federal Income Tax will be withheld at a rate of 20% for everyone. FICA Social Security rate is 6.2% and Medicare is 1.45%. When you called your State, they told you that your State Unemployment tax rate would be 5.4%. Recall that the Federal Unemployment rate is the same for all businesses. Required: Complete the Payroll Report worksheet Use the Transactions worksheet to complete your Journal Required: Complete the payroll report below using the information provided on the Information tab. January Employee Name Hours worked Hourly rate Gross Pay Federal Withholding Social Security Medicare Net Required: Use the transactions below to complete your Journal on the next worksheet. Tip: Refer to pages 473-476 of your text for examples. Transactions January 1 Record accrued payroll for January using the Payroll Report 2 Record employer payroll taxes for the same Payroll Report (hint - remember to record FUTA and SUTA as well) 7 Record payment of cash to employees. Required: Enter the journal entries from the transactions found on the Transaction worksheet Journal Entries Date Account Title Debit Credit
Answer #1
Employee
Name
Hours
worked
Hourly
rate
Gross
pay
Federal
withholding
Social
security
Medicare Net
1 2 3=1*2 4=3*20% 5=3*(6.2%) 6=3*(1.45%) 7=3-4-5-6
John Doe 40 25 1000 200 62 14.5 723.5
Jane Jones 40 24 960 192 59.52 13.92 694.56
Jessie Smith 40 17 680 136 42.16 9.86 491.98
Erik Ackers 25 15 375 75 23.25 5.44 271.31
Nan Schmit 30 17 510 102 31.62 7.40 368.99
Larry Johnson 35 18 630 126 39.06 9.14 455.81
Christopher Hay 40 18 720 144 44.64 10.44 520.92
Total 4875 975 302.25 70.69 3527.06
Journal entry to record accrued salaries:
Date Account title Debit Credit
Salaries and wages expense 4875
Federal taxes payable 975
Social security taxes payable 302.25
Medicare taxes payable 70.69
Salaries and wages payable 3527.06
Employer payroll taxes:
Employee
Name
Gross
pay
Social
security
Medicare State
Unemployment
tax rate
Federal
Unemployment
tax rate
1 2=1*(6.2%) 3=1*(1.45%) 4=1*(5.4%) 5=1*(0.6%)
John Doe 1000 62 14.5 54 6
Jane Jones 960 59.52 13.92 51.84 5.76
Jessie Smith 680 42.16 9.86 36.72 4.08
Erik Ackers 375 23.25 5.4375 20.25 2.25
Nan Schmit 510 31.62 7.395 27.54 3.06
Larry Johnson 630 39.06 9.135 34.02 3.78
Christopher Hay 720 44.64 10.44 38.88 4.32
Total 4875 302.25 70.6875 263.25 29.25
Date Account title Debit Credit
Payroll tax expense 665.44
Social security taxes payable 302.25
Medicare taxes payable 70.69
State unemplyment taxes payable 263.25
Federeal unemplyment taxes payable 29.25
(To record employer payroll tax expenses)
Salaries and wages payable 3527.06
Cash 3527.06
(Cash paid to employees)